Suppose a1 = (1, 0, 1), a2 = (0, 1, -2) and
a3 = (-1, -1, 0) are vectors in R³ and
f : R³ -> R is a linear functional such that
f(a1) = 1, f(a2) = -1 and f(a3) = 3. If
a = (p,q,r) belongs to R3, find f(a).
1
Expert's answer
2020-11-16T20:29:06-0500

A mapping "f:V\to \Bbb{F}" from the vector space "V" to scalar field "\Bbb{F}" is called a linear functional iff:


"f(\alpha x+\beta y)=\alpha f(x)+\beta f(y)"

holds for all "x,y" in "V" and for all "\alpha, \beta" in "\Bbb{F}"

Given


"a=(p,q,r)=\lambda_1(1,0,1)+\lambda_2(0,1,-2)+\lambda_3(-1,-1,0)"

"\begin{matrix}
 \lambda_1-\lambda_3=p \\
 \lambda_2-\lambda_3=q \\
 \lambda_1-2\lambda_2=r \\
\end{matrix}"

"\begin{matrix}
 \lambda_1=\lambda_3+p \\
 \lambda_2=\lambda_3+q \\
 \lambda_3+p-2\lambda_3-2q=r \\
\end{matrix}"


"\begin{matrix}
 \lambda_1=2p-2q-r \\
 \lambda_2=p-q-r \\
 \lambda_3=p-2q-r \\
\end{matrix}"

"a=(2p-2q-r)a_1+(p-q-r)a_2+(p-2q-r)a_3"

"f(a)=(2p-2q-r)f(a_1)+(p-q-r)f(a_2)+"

"+(p-2q-r)f(a_3)="

"=(2p-2q-r)(1)+(p-q-r)(-1)+"

"+(p-2q-r)(3)="

"=2p-2q-r-p+q+r+3p-6q-3r="

"=4p-7q-3r"

"f(a)=4p-7q-3r"